OpenStudy (anonymous):

How many definite integrals would be required to represent the area of the region enclosed by the curves y = cos2x sinx and y = 0.03x2, assuming you could not use the absolute value function? 1 2 3 4 5 Find the area of the region bounded by the curve y = f(x) = x3 – 4x + 1 and the tangent line to the curve y = f(x) at (–1,4). You get: 8.25 6.25

OpenStudy (btaylor):

For the first one, look at the graph (attached). Since you can't use the absolute value functions, how many different regions are defined?

OpenStudy (anonymous):

4?

OpenStudy (btaylor):

Actually 5. There is a small area (I don't know if you can even see the shading) that you probably missed. For the second one, first you must find the equation of the tangent line first, the integrate.

OpenStudy (anonymous):

thnxa

OpenStudy (anonymous):

does that mean find derivative

OpenStudy (btaylor):

Yes, then evaluate it at the given x-value.

OpenStudy (anonymous):

could u help me with some more plz

OpenStudy (btaylor):

sure.

OpenStudy (anonymous):

thnxs

OpenStudy (btaylor):

It looks like you are dealing with average value. You can find this with: \[\frac{1}{b-a}\int\limits_{a}^{b}f(x)dx\] I assume you can use a calculator?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@BTaylor

OpenStudy (btaylor):

Ok, so can you set up the integral for #6?

OpenStudy (anonymous):

ok give me a sec and check my work when I post k? and thnxs

OpenStudy (btaylor):

sure thing.

OpenStudy (anonymous):

hold one

OpenStudy (anonymous):

\[1/c-1\int\limits_{1}^{c}(9\pi/x^2)\cos(\pi/x)dx\]

OpenStudy (anonymous):

is this right?

OpenStudy (btaylor):

yes, but according to the problem it is set equal to -.9 Now, multiply both sides by (c-1). Using your calculator, find the indefinite integral.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

\[(1/c-1)\int\limits_{1}^{c}(9\pi/x^2)\cos(\pi/x)=-0.9\]

OpenStudy (anonymous):

I need some help @BTaylor

OpenStudy (anonymous):

@sourwing could u help me plz

OpenStudy (btaylor):

sorry, had to go for a sec. You will end up with \[\int\limits_{1}^{c}\frac{9 \pi}{x^2}\cos \left( \frac{\pi}{x} \right)dx=-0.9(c-1)\]

OpenStudy (anonymous):

then what?

OpenStudy (btaylor):

take the antiderivative/indefinite integral of the left side. Using the fundamental theorem of calculus, set it up so you have F(c)-F(1) -> F(x) is the antiderivative of f(x) evaluated at x.

OpenStudy (btaylor):

From there you should be able to solve.

OpenStudy (anonymous):

could u show me how to do it on this one plz

OpenStudy (btaylor):

I'll try. The indefinite integral is actually quite simple: \(F(x)=-9sin \left(\frac{\pi}{x} \right)+C \)

OpenStudy (btaylor):

So, we get: \[-9\sin \left( \frac{\pi}{c} \right) +9\sin(1)=-.9(c-1)\]

OpenStudy (btaylor):

Use your calculator to solve for c.

OpenStudy (btaylor):

wait...looking at the question, just plug in for c the different multiple choice values. Saves you work to solve. *facepalm*

OpenStudy (anonymous):

great thnxs, could u help me with another one plzzz

OpenStudy (btaylor):

sure? which one?

OpenStudy (anonymous):

7

OpenStudy (btaylor):

This one is slightly easier. You know you need to find the average value, and this time it gives you the bounds. Using the formula \[average ~ value = \frac{1}{b-a}\int\limits_{a}^{b}f(x)dx\]can you find the average value of that polynomial?

OpenStudy (anonymous):

yah, but im still working on the other one so brb

OpenStudy (anonymous):

is 6 e?@BTaylor

OpenStudy (anonymous):

@BTaylor

OpenStudy (anonymous):

is 6 e?

OpenStudy (btaylor):

Sounds right to me.

OpenStudy (anonymous):

cool ok now 7.

OpenStudy (anonymous):

can u help me with 8 plz

OpenStudy (anonymous):

@BTaylor

OpenStudy (btaylor):

just use that average value formula and plug it into your calculator.

OpenStudy (anonymous):

is that the answer?

OpenStudy (btaylor):

yep. Since the question (for #8) asks for the average value, you just need to answer with the average value.

OpenStudy (anonymous):